# CLIP-AGIQA: Boosting the Performance of AI-Generated Image Quality Assessment with CLIP

### v2

@@ Abstract
With the rapid development of generative technologies, AI-Generated Images (AIGIs) have been widely applied in various aspects of daily life. However, due to the immaturity of the technology, the quality of the generated images varies, so it is important to develop quality assessment techniques for the generated images. Although some models have been proposed to assess the quality of generated images, they are inadequate when faced with the ever-increasing and diverse categories of generated images. Consequently, the development of more advanced and effective models for evaluating the quality of generated images is urgently needed. Recent research has explored the significant potential of the visual language model CLIP in image quality assessment, finding that it performs well in evaluating the quality of natural images. However, its application to generated images has not been thoroughly investigated. In this paper, we build on this idea and further explore the potential of CLIP in evaluating the quality of generated images. We design CLIP-AGIQA, a CLIP-based regression model for quality assessment of generated images, leveraging rich visual and textual knowledge encapsulated in CLIP. Particularly, we implement multi-category learnable prompts to fully utilize the textual knowledge in CLIP for quality assessment. Extensive experiments on several generated image quality assessment benchmarks, including AGIQA-3K and AIGCIQA2023, demonstrate that CLIP-AGIQA outperforms existing IQA models, achieving excellent results in evaluating the quality of generated images. Code is available at https://github.com/wzczc/CLIP-AGIQA.
@@
reality, generative technologies has become an integral part of our daily experiences. However, alongside these technological advancements, assessing the quality of generated images has become an emerging issue. Due to the immaturity of the technology, the quality of generated images is uneven, which can lead to unsatisfactory user experiences in some applications [14]. Therefore, developing techniques to effectively evaluate the quality of generated images is particularly important.
Quality assessment of generated images involves evaluating various dimensions through subjective and objective methods, such as the perceptual quality and the content accuracy with respect to input prompts. Recent efforts have focused on creating comprehensive databases for subjective quality assessment based on human perception and developing approaches to enhance evaluation performance [14, 29, 12]. Despite these advancements, existing methods struggle to keep pace with the increasing diversity of generated images. For instance, in the field of text-to-image (T2I) generative models alone, there have been at least 20 representative T2I AGI models up to 2023, as indicated by recent statistics [2, 34]. Therefore, more research is needed to meet the quality assessment demands in this field.
Recent research has begun to explore CLIP’s [18] (Contrastive Language-Image Pre-training) potential in image quality assessment, revealing its effectiveness in evaluating natural images [24]. CLIP demonstrates strong performance across various visual and multimodal tasks due to its extensive pre-training on language-image data. However, since CLIP is pre-trained on natural images, it may have problems to model the quality distribution of generated images effectively, leaving a gap in this area. To address this, we propose CLIP-AGIQA, a CLIP-based regression model that leverages CLIP’s comprehensive visual and textual knowledge to evaluate the quality of generated images. First, we design various prompts representing different quality levels to input into CLIP’s text encoder, mitigating semantic ambiguities. Second, by introducing a learnable prompts strategy and utilizing multiple quality-related auxiliary prompts, we make full use of CLIP’s textual knowledge. Last, our regression network then maps CLIP features to quality scores, effectively adapting CLIP’s capabilities to the task of generated image quality assessment, thereby enhancing the model’s performance. The specific performance of our CLIP-AGIQA can be seen in Fig.1.
@@ 2.1 Image Quality Assessment
Traditional image quality assessment aims to evaluate the quality of natural images, including aspects like noise, blur, compression artifacts, etc [3]. It is categorized into three types: full-reference, reduced-reference, and no-reference. Full-reference methods compare the original and test images, commonly using metrics like PSNR and SSIM [26]. Reduced-reference methods utilize partial information from a reference image, such as RRED [21] and OSVP [27]. No-reference methods directly assess image quality using machine learning and deep learning techniques, such as BRISQUE [16], IQA-CNN [9] and RankIQA [15].
In recent years, with the development of generative technologies, assessing the quality of generated images has become increasingly important. Due to potential abnormal distortions or unrealistic structures in generated images, evaluation focuses on visual perception, including authenticity, naturalness, and coherence. Common metrics include Inception Score (IS) for assessing image quality and diversity based on classification results and KL divergence [19], Fréchet Inception Distance (FID) for evaluating visual quality by comparing feature distributions of real and generated images [7], and CLIP Score, which assesses image quality based on similarity between generated images and textual descriptions [6].
Recently, datasets like AGIQA-3K [14] and PKU-I2IQA [33] have been proposed to facilitate benchmark experiments for IQA models, focusing on the quality assessment of generated images. AGIQA-3K provides a comprehensive and diverse subjective quality database covering various generated images from GAN, autoregressive, and diffusion models. PKU-I2IQA, the first image-to-image AIGC quality assessment database based on human perception, also conducts benchmark experiments on different IQA models. Additionally, models such as ImageReward [29] and HPS [28] construct datasets for generated images from the perspective of human preferences and proposed corresponding evaluation models, providing a benchmark for quality assessment in terms of human preferences for generated images. Despite these advancements, there remains a scarcity of specialized models for assessing the quality of generated images, necessitating further research to advance this field.
@@ 2.2 CLIP-Based Methods
CLIP [18] is a large-scale vision-language pretrained model that leverages contrastive learning to achieve cross-modal knowledge understanding. It has demonstrated strong transfer capabilities across various visual tasks such as semantic segmentation (LSeg [13]), object detection (ViLD [4]), and image generation (CLIPasso [23]).
CLIP-IQA [24] is the first work to explore CLIP in image quality assessment tasks, demonstrating that CLIP can be effectively extended to image quality evaluation. Due to the significant impact of linguistic ambiguity in quality assessment tasks [11], phrases such as "a rich image" can be particularly problematic. This phrase could either refer to an image with rich content or an image associated with wealth. CLIP-IQA design an antonym prompt strategy to leverage CLIP’s prior knowledge. However, due to the limited variety of prompts, this approach can result in inaccurate quality predictions. Moreover, this work only explored the performance of CLIP in natural image quality assessment tasks and did not address generated images. Building on this idea, we further investigate the performance of CLIP in evaluating the quality of generated images and propose a CLIP-based quality assessment regression model. By simultaneously fine-tuning our designed multi-class learnable prompts and the regression network added after CLIP, we achieve superior performance in assessing the quality of generated images.
Notably, recent methods [8, 10, 36] also explore CLIP for IQA, with many focusing on aesthetic evaluation. These methods stand out for their pioneering efforts in multi-modality integration for low-level vision and their impressive zero-shot performance. However, since CLIP is pre-trained on natural image-text pairs, directly using CLIP in a zero-shot manner to evaluate the quality of generated images, as done in the aforementioned methods, does not yield optimal results. Therefore, we train a CLIP-based model using generated images to better model the quality distribution of generated images.
@@ 3.1 Preliminary on IQA Models
In CLIP-IQA [24], only the visual encoder is used to extract visual features, and then an antonym prompt strategy is employed to compute the cosine similarity with the visual features to predict the quality score. Specifically, CLIP-IQA adopts antonym prompts (e.g., "Good photo." and "Bad photo.") as a pair for each prediction. Let represent the features from the image, and and be the features from the two prompts with opposite meanings. The cosine similarity is computed as follows [24]:
@@
When a pair of adjectives is used, the ambiguity of one prompt is reduced by its antonym, casting the task as a binary classification where the final score is regarded as a relative similarity [24]. Although this method effectively leverages the prior knowledge of CLIP, the predicted quality score is solely dependent on the contrastive similarity, which is not accurate. Therefore, in our design, we improve the network by using a regression model to predict the quality score, enhancing the precision of the prediction and better adapting CLIP to the quality assessment task after further reducing ambiguity with more fine-grained quality-related adjectives.
@@ 3.2 Overview of CLIP-AGIQA
The overall framework of our method is shown in Fig.2. CLIP-AGIQA consists of four components: learnable context, quality category, image quality regression, and the text encoder and image encoder in CLIP. In addition to the regression design, to better utilize the prior knowledge of the CLIP model, we incorporate learnable context for fine-tuning, inspired by the CoOp approach [37]. We also introduce additional quality category to address the ambiguity issues mentioned in CLIP-IQA. These two types of text-related information together form supplementary textual information to assist CLIP in adapting to the task of generative image quality assessment.
@@
Since prompt engineering is a significant challenge in the application of CLIP, and the design of prompts can greatly impact performance, even with extensive manual tuning, the resulting prompts are by no means guaranteed to be optimal for downstream tasks [37]. Therefore, we abandon traditional subjective prompt settings in favor of a learnable prompt strategy. CLIP is sensitive to the choice of prompts, so we need to design a suitable set to leverage its prior knowledge. Similar to CoOp [37], we avoid manual prompt adjustments by modeling the context words using continuous vectors, which are end-to-end learned from the data, while freezing a large number of CLIP’s pre-trained parameters. Specifically, as shown in Fig.2, we use learnable context. We employ a unified context version from CoOp, where all prompts share the same context. The prompt design for the text encoder is as follows:
@@
We utilize the text encoder and image encoder from CLIP. The text encoder is based on a Transformer architecture [22] and is responsible for generating text representations from natural language. In contrast, the image encoder is designed to map high-dimensional images into a low-dimensional embedding space. This encoder’s architecture can resemble a CNN like ResNet-50 [5] or a Vision Transformer (ViT) [1]. In our setup, we employ these encoders separately to process our input textual information and image information , generating intermediate features used to predict quality score.
@@ 4.1 Experimental Settings
To validate the effectiveness of our method, we conduct evaluations on two quality assessment benchmarks for generated images: AGIQA-3K [14] ans AIGCIQA2023 [25]. AGIQA-3K is a database containing 2,982 AI-generated images produced by six different models, including GAN-based, auto-regression-based, and diffusion-based models and subjective experiments are organized to obtain MOS (Mean Opinion Score) labels in terms of perceptual quality, which range from 0 to 5. AIGCIQA2023 collects over 2000 images using 100 prompts and six state-of-the-art text-to-image generation models, and quality and authenticity ratings are obtained by subjective experiments, which are ultimately scaled to a range of 0-100.
@@
The proposed CLIP-AGIQA is implemented in PyTorch and trained on 1 NVIDIA A100 GPU. ViT-B/16 [1] is used as the image encoder’s backbone, and SGD is applied to optimize the network with an initial learning rate of 0.002. The training process was conducted over 100 epochs with a batch size of 32 and a learnable context length of 16. For learning rate scheduling, we employed a cosine annealing strategy, allowing the learning rate to decrease gradually throughout the training. Additionally, we implemented a warm-up phase during the first epoch, where the learning rate was held constant at .
@@ 4.2 Experiment on Different Datasets
We focus on exploring the potential of CLIP-AGIQA in overall quality perception assessment. We conduct experiments on two widely used AGIQA benchmarks: AGIQA-3K [14] ans AIGCIQA2023 [25]. We also compare CLIP-AGIQA with different IQA methods, including handcrafted-based methods such as CEIQ [32], NIQE [17] and BRISQUE [16], and several learning-based methods like DBCNN [35], CLIP-IQA [24] and CNNIQA [9].
@@
| FID [7] | 0.1860 | 0.1733 | 0.1158 |
| CEIQ [32] | 0.4166 | 0.3228 | 0.2220 |
| NIQE [17] | 0.5171 | 0.5623 | 0.3876 |
| GMLF [31] | 0.8181 | 0.6987 | 0.5119 |
| CNNIQA [9] | 0.8469 | 0.7478 | 0.5580 |
| DBCNN [35] | 0.8759 | 0.8207 | 0.6336 |
| CLIP-IQA [24] | 0.8053 | 0.8426 | 0.6468 |
@@
| NIQE [17] | 0.5218 | 0.5060 | 0.3420 | 3954 | 3715 | 2453 |
| BRISQUE [16] | 0.6389 | 0.6239 | 0.4291 | 0.4796 | 0.4705 | 0.3142 |
| HOSA [30] | 0.6561 | 0.6317 | 0.4311 | 0.4985 | 0.4716 | 0.3101 |
| CNNIQA [9] | 0.7937 | 0.7160 | 0.4955 | 0.5734 | 0.5958 | 0.4085 |
| Resnet18 [5] | 0.7763 | 0.7583 | 0.5360 | 0.6528 | 0.6701 | 0.4740 |
| VGG16 [20] | 0.7973 | 0.7961 | 0.5843 | 0.6807 | 0.6660 | 0.4813 |
| VGG19 [20] | 0.8402 | 0.7733 | 0.5376 | 0.6565 | 0.6674 | 0.4843 |

### v3

@@ Abstract
With the rapid development of generative technologies, AI-Generated Images (AIGIs) have been widely applied in various aspects of daily life. However, due to the immaturity of the technology, the quality of the generated images varies, so it is important to develop quality assessment techniques for the generated images. Although some models have been proposed to assess the quality of generated images, they are inadequate when faced with the ever-increasing and diverse categories of generated images. Consequently, the development of more advanced and effective models for evaluating the quality of generated images is urgently needed. Recent research has explored the significant potential of the visual language model CLIP in image quality assessment, finding that it performs well in evaluating the quality of natural images. However, its application to generated images has not been thoroughly investigated. In this paper, we build on this idea and further explore the potential of CLIP in evaluating the quality of generated images. We design CLIP-AGIQA, a CLIP-based regression model for quality assessment of generated images, leveraging rich visual and textual knowledge encapsulated in CLIP. Particularly, we implement multi-category learnable prompts to fully utilize the textual knowledge in CLIP for quality assessment. Extensive experiments on several generated image quality assessment benchmarks, including AGIQA-3K and AIGCIQA2023, demonstrate that CLIP-AGIQA outperforms existing IQA models, achieving excellent results in evaluating the quality of generated images.
@@
reality, generative technologies has become an integral part of our daily experiences. However, alongside these technological advancements, assessing the quality of generated images has become an emerging issue. Due to the immaturity of the technology, the quality of generated images is uneven, which can lead to unsatisfactory user experiences in some applications [li2023agiqa]. Therefore, developing techniques to effectively evaluate the quality of generated images is particularly important.
Quality assessment of generated images involves evaluating various dimensions through subjective and objective methods, such as the perceptual quality and the content accuracy with respect to input prompts. Recent efforts have focused on creating comprehensive databases for subjective quality assessment based on human perception and developing approaches to enhance evaluation performance [li2023agiqa, xu2024imagereward, kirstain2023pick]. Despite these advancements, existing methods struggle to keep pace with the increasing diversity of generated images. For instance, in the field of text-to-image (T2I) generative models alone, there have been at least 20 representative T2I AGI models up to 2023, as indicated by recent statistics [frolov2021adversarial, zhang2023text]. Therefore, more research is needed to meet the quality assessment demands in this field.
Recent research has begun to explore CLIP’s [radford2021learning] (Contrastive Language-Image Pre-training) potential in image quality assessment, revealing its effectiveness in evaluating natural images [wang2023exploring]. CLIP demonstrates strong performance across various visual and multimodal tasks due to its extensive pre-training on language-image data. However, since CLIP is pre-trained on natural images, it may have problems to model the quality distribution of generated images effectively, leaving a gap in this area. To address this, we propose CLIP-AGIQA, a CLIP-based regression model that leverages CLIP’s comprehensive visual and textual knowledge to evaluate the quality of generated images. First, we design various prompts representing different quality levels to input into CLIP’s text encoder, mitigating semantic ambiguities. Second, by introducing a learnable prompts strategy and utilizing multiple quality-related auxiliary prompts, we make full use of CLIP’s textual knowledge. Last, our regression network then maps CLIP features to quality scores, effectively adapting CLIP’s capabilities to the task of generated image quality assessment, thereby enhancing the model’s performance. The specific performance of our CLIP-AGIQA can be seen in Fig.1.
@@ 2.1 Image Quality Assessment
Traditional image quality assessment aims to evaluate the quality of natural images, including aspects like noise, blur, compression artifacts, etc [gu2020giqa]. It is categorized into three types: full-reference, reduced-reference, and no-reference. Full-reference methods compare the original and test images, commonly using metrics like PSNR and SSIM [wang2004image]. Reduced-reference methods utilize partial information from a reference image, such as RRED [soundararajan2011rred] and OSVP [wu2016orientation]. No-reference methods directly assess image quality using machine learning and deep learning techniques, such as BRISQUE [mittal2012no], IQA-CNN [kang2014convolutional] and RankIQA [liu2017rankiqa].
In recent years, with the development of generative technologies, assessing the quality of generated images has become increasingly important. Due to potential abnormal distortions or unrealistic structures in generated images, evaluation focuses on visual perception, including authenticity, naturalness, and coherence. Common metrics include Inception Score (IS) for assessing image quality and diversity based on classification results and KL divergence [salimans2016improved], Fréchet Inception Distance (FID) for evaluating visual quality by comparing feature distributions of real and generated images [heusel2017gans], and CLIP Score, which assesses image quality based on similarity between generated images and textual descriptions [hessel2021clipscore].
Recently, datasets like AGIQA-3K [li2023agiqa] and PKU-I2IQA [yuan2023pku] have been proposed to facilitate benchmark experiments for IQA models, focusing on the quality assessment of generated images. AGIQA-3K provides a comprehensive and diverse subjective quality database covering various generated images from GAN, autoregressive, and diffusion models. PKU-I2IQA, the first image-to-image AIGC quality assessment database based on human perception, also conducts benchmark experiments on different IQA models. Additionally, models such as ImageReward [xu2024imagereward] and HPS [wu2023human] construct datasets for generated images from the perspective of human preferences and proposed corresponding evaluation models, providing a benchmark for quality assessment in terms of human preferences for generated images. Despite these advancements, there remains a scarcity of specialized models for assessing the quality of generated images, necessitating further research to advance this field.
@@ 2.2 CLIP-Based Methods
CLIP [radford2021learning] is a large-scale vision-language pretrained model that leverages contrastive learning to achieve cross-modal knowledge understanding. It has demonstrated strong transfer capabilities across various visual tasks such as semantic segmentation (LSeg [li2022language]), object detection (ViLD [gu2021open]), and image generation (CLIPasso [vinker2022clipasso]).
CLIP-IQA [wang2023exploring] is the first work to explore CLIP in image quality assessment tasks, demonstrating that CLIP can be effectively extended to image quality evaluation. Due to the significant impact of linguistic ambiguity in quality assessment tasks [khurana2023natural], phrases such as "a rich image" can be particularly problematic. This phrase could either refer to an image with rich content or an image associated with wealth. CLIP-IQA design an antonym prompt strategy to leverage CLIP’s prior knowledge. However, due to the limited variety of prompts, this approach can result in inaccurate quality predictions. Moreover, this work only explored the performance of CLIP in natural image quality assessment tasks and did not address generated images. Building on this idea, we further investigate the performance of CLIP in evaluating the quality of generated images and propose a CLIP-based quality assessment regression model. By simultaneously fine-tuning our designed multi-class learnable prompts and the regression network added after CLIP, we achieve superior performance in assessing the quality of generated images.
Notably, recent methods [hou2023towards, ke2023vila, zhang2023blind] also explore CLIP for IQA, with many focusing on aesthetic evaluation. These methods stand out for their pioneering efforts in multi-modality integration for low-level vision and their impressive zero-shot performance. However, since CLIP is pre-trained on natural image-text pairs, directly using CLIP in a zero-shot manner to evaluate the quality of generated images, as done in the aforementioned methods, does not yield optimal results. Therefore, we train a CLIP-based model using generated images to better model the quality distribution of generated images.
@@ 3.1 Preliminary on IQA Models
In CLIP-IQA [wang2023exploring], only the visual encoder is used to extract visual features, and then an antonym prompt strategy is employed to compute the cosine similarity with the visual features to predict the quality score. Specifically, CLIP-IQA adopts antonym prompts (e.g., "Good photo." and "Bad photo.") as a pair for each prediction. Let represent the features from the image, and and be the features from the two prompts with opposite meanings. The cosine similarity is computed as follows [wang2023exploring]:
@@
When a pair of adjectives is used, the ambiguity of one prompt is reduced by its antonym, casting the task as a binary classification where the final score is regarded as a relative similarity [wang2023exploring]. Although this method effectively leverages the prior knowledge of CLIP, the predicted quality score is solely dependent on the contrastive similarity, which is not accurate. Therefore, in our design, we improve the network by using a regression model to predict the quality score, enhancing the precision of the prediction and better adapting CLIP to the quality assessment task after further reducing ambiguity with more fine-grained quality-related adjectives.
@@ 3.2 Overview of CLIP-AGIQA
The overall framework of our method is shown in Fig.2. CLIP-AGIQA consists of four components: learnable context, quality category, image quality regression, and the text encoder and image encoder in CLIP. In addition to the regression design, to better utilize the prior knowledge of the CLIP model, we incorporate learnable context for fine-tuning, inspired by the CoOp approach [zhou2022learning]. We also introduce additional quality category to address the ambiguity issues mentioned in CLIP-IQA. These two types of text-related information together form supplementary textual information to assist CLIP in adapting to the task of generative image quality assessment.
@@
Since prompt engineering is a significant challenge in the application of CLIP, and the design of prompts can greatly impact performance, even with extensive manual tuning, the resulting prompts are by no means guaranteed to be optimal for downstream tasks [zhou2022learning]. Therefore, we abandon traditional subjective prompt settings in favor of a learnable prompt strategy. CLIP is sensitive to the choice of prompts, so we need to design a suitable set to leverage its prior knowledge. Similar to CoOp [zhou2022learning], we avoid manual prompt adjustments by modeling the context words using continuous vectors, which are end-to-end learned from the data, while freezing a large number of CLIP’s pre-trained parameters. Specifically, as shown in Fig.2, we use learnable context. We employ a unified context version from CoOp, where all prompts share the same context. The prompt design for the text encoder is as follows:
@@
We utilize the text encoder and image encoder from CLIP. The text encoder is based on a Transformer architecture [vaswani2017attention] and is responsible for generating text representations from natural language. In contrast, the image encoder is designed to map high-dimensional images into a low-dimensional embedding space. This encoder’s architecture can resemble a CNN like ResNet-50 [he2016deep] or a Vision Transformer (ViT) [dosovitskiy2020image]. In our setup, we employ these encoders separately to process our input textual information and image information , generating intermediate features used to predict quality score.
@@ 4.1 Experimental Settings
To validate the effectiveness of our method, we conduct evaluations on two quality assessment benchmarks for generated images: AGIQA-3K [li2023agiqa] ans AIGCIQA2023 [wang2023aigciqa2023]. AGIQA-3K is a database containing 2,982 AI-generated images produced by six different models, including GAN-based, auto-regression-based, and diffusion-based models and subjective experiments are organized to obtain MOS (Mean Opinion Score) labels in terms of perceptual quality, which range from 0 to 5. AIGCIQA2023 collects over 2000 images using 100 prompts and six state-of-the-art text-to-image generation models, and quality and authenticity ratings are obtained by subjective experiments, which are ultimately scaled to a range of 0-100.
@@
The proposed CLIP-AGIQA is implemented in PyTorch and trained on 1 NVIDIA A100 GPU. ViT-B/16 [dosovitskiy2020image] is used as the image encoder’s backbone, and SGD is applied to optimize the network with an initial learning rate of 0.002. The training process was conducted over 100 epochs with a batch size of 32 and a learnable context length of 16. For learning rate scheduling, we employed a cosine annealing strategy, allowing the learning rate to decrease gradually throughout the training. Additionally, we implemented a warm-up phase during the first epoch, where the learning rate was held constant at .
@@ 4.2 Experiment on Different Datasets
We focus on exploring the potential of CLIP-AGIQA in overall quality perception assessment. We conduct experiments on two widely used AGIQA benchmarks: AGIQA-3K [li2023agiqa] ans AIGCIQA2023 [wang2023aigciqa2023]. We also compare CLIP-AGIQA with different IQA methods, including handcrafted-based methods such as CEIQ [yan2019no], NIQE [mittal2012making] and BRISQUE [mittal2012no], and several learning-based methods like DBCNN [zhang2018blind], CLIP-IQA [wang2023exploring] and CNNIQA [kang2014convolutional].
@@
| FID [heusel2017gans] | 0.1860 | 0.1733 | 0.1158 |
| CEIQ [yan2019no] | 0.4166 | 0.3228 | 0.2220 |
| NIQE [mittal2012making] | 0.5171 | 0.5623 | 0.3876 |
| GMLF [xue2014blind] | 0.8181 | 0.6987 | 0.5119 |
| CNNIQA [kang2014convolutional] | 0.8469 | 0.7478 | 0.5580 |
| DBCNN [zhang2018blind] | 0.8759 | 0.8207 | 0.6336 |
| CLIP-IQA [wang2023exploring] | 0.8053 | 0.8426 | 0.6468 |
@@
| NIQE [mittal2012making] | 0.5218 | 0.5060 | 0.3420 | 0.3954 | 0.3715 | 0.2453 |
| BRISQUE [mittal2012no] | 0.6389 | 0.6239 | 0.4291 | 0.4796 | 0.4705 | 0.3142 |
| HOSA [xu2016blind] | 0.6561 | 0.6317 | 0.4311 | 0.4985 | 0.4716 | 0.3101 |
| CNNIQA [kang2014convolutional] | 0.7937 | 0.7160 | 0.4955 | 0.5734 | 0.5958 | 0.4085 |
| Resnet18 [he2016deep] | 0.7763 | 0.7583 | 0.5360 | 0.6528 | 0.6701 | 0.4740 |
| VGG16 [simonyan2014very] | 0.7973 | 0.7961 | 0.5843 | 0.6807 | 0.6660 | 0.4813 |
| VGG19 [simonyan2014very] | 0.8402 | 0.7733 | 0.5376 | 0.6565 | 0.6674 | 0.4843 |